# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 6. Proof of main results
For any test function $\phi\in\mathcal{D}_{T}$ and $t_{0}\in(T_{*},T)$, we have
@@
for all $p<\infty$. Furthermore, $u$ lies in the Koch–Tataru path space $X_{[0,T-T_{*}]}$ and is weak-* continuous in time into $BMO^{-1}$. When $d\geq 3$, we additionally have
@@
First we address the boundedness in $\dot{W}^{-1,\infty}$. Recall that, by definition, $u=U+v^{N_{0}}+w^{N_{0}}$. Due to the criticality of the norm, we may equivalently estimate $w$ which, by construction, lies in $B_{X}(0,\delta)$. From this, (4.18), and (4.16), it follows
$\|w(t)\|_{\dot{W}^{-1,\infty}}$ $\lesssim\int_{0}^{t}\|w\otimes w+2(v+U^{1/N_{0}})\odot w+f\|_{C^{\varepsilon}}$
$\lesssim\int_{0}^{t}(s^{-1+\alpha-\varepsilon}+1)ds$
@@
$\|R_{k}(t)\|_{\infty}$ $\lesssim\int_{0}^{t}\|\bar{v}_{k+1}\otimes\bar{v}_{k+1}\|_{C^{\varepsilon}}\lesssim\int_{0}^{t}\sum_{j}N_{j,k+1}^{2+\varepsilon}e^{-2N_{j,k+1}^{2}t}ds\lesssim tN_{J_{d},k+1}^{2+\varepsilon}.$
@@
Next, we address the weak-* continuity in time into $BMO^{-1}$. For every $t\in[-T_{*},T-T_{*}]\setminus\{0\}$, $u$ is a classical solution and continuity is obvious. At $t=0$, left continuity is clear because $u$ agrees with $U$ on $T_{*},0]$. (Note that left continuity is vacuous if $T_{*}=0$.) To show right continuity at $t=0$, it suffices to show $u|_{[0,T-T_{*}]}\in X_{[0,T-T_{*}]}$ due to the recent result [hou2024regularity] of Hedong Hou. To this end, we decompose
@@
$\sup_{t\in(0,1-T_{*}]}t^{\frac{1}{2}-\epsilon}\|f(t)\|_{L^{\infty}}<\infty\implies f\in X_{1-T_{*}}$ (6.1)
@@
$e^{t\Delta}N_{j,k}\Delta\psi_{j,k}+\bar{v}_{j,k}$ $=N_{j,k}\Delta(e^{t\Delta}-e^{N_{j,k}^{2}t})\psi_{j,k}$
@@
That $u_{2}\in X_{1-T_{*}}$ is immediate from Proposition 4.2 and (6.1).
@@
where the supremum is taken over cubes $Q\subset\mathbb{R}^{d}$; clearly, by periodicity, it suffices to consider only cubes below side length $1$, say. We proceed as in [CoiculescuPalasek2025]*Proposition 3.8. Let $k(Q)$ and $j(Q)$ be such that
@@
Thus, the upper bound $\|u(t)\|_{L^{\infty}}\lesssim\frac{1}{\sqrt{t}}$ on $[0,T)$. To obtain the lower bound, we note there exists a time sequence $t_{n}\to 0+$ such that the principal velocity field $v$ satisfies
@@
We claim that the family of solutions $u^{(\sigma_{n})}$ converges to $u^{(0)}$ in $L_{t}^{\infty}([0,T];BMO^{-1}_{x}(\mathbb{R}^{d}))$ with the weak-* $BMO^{-1}$ topology. Indeed, by the rescaling, $\hat{v}^{C/\sigma_{n}}+\hat{w}^{C/\sigma_{n}}$ is supported only in $(C/\sigma_{n})\mathbb{Z}^{d}$. Since they are zero average the support must be contained in $\{\xi\in\mathbb{Z}^{d}:|\xi|>C/\sigma_{n}\}$. Further, by criticality, we have
@@
Having shown inclusion in the Koch–Tataru space on the lifetime of the solution, the uniform bound in $BMO^{-1}$ follows from the theorem of [hou2024regularity]. Thus we conclude that for any test function $\varphi$ in the predual, we have
@@
$\lesssim\|P_{>C/(2\sigma_{n})}\varphi\|_{L_{t}^{1}H^{1,1}}\|u^{(\sigma_{n})}-U\|_{L_{t}^{\infty}BMO^{-1}}.$
@@
$\|\nabla^{n}e^{t\Delta}\mathcal{R}\mathbb{P}(ae^{i\xi\cdot x})\|_{C^{\kappa}}$ $\lesssim|\xi|^{n+\kappa}\left(|\xi|^{-1}\sum_{i=0}^{m-1}A_{i,\kappa}e^{-|\xi|^{2}/4}+\sum_{i=m}^{n+2m}A_{i,\kappa}\right).$
@@
The Leray projection is well-known to be bounded on $C^{\kappa}$ so it can be removed immediately. For the rest, we recall the following identity from [DeLellisSzekelyhidi2013]*Prop. 5.2:
@@
Since we can write $\mathcal{R}f=\frac{1}{2}\Delta^{-1}(\nabla\odot\mathbb{P}f+3\nabla\odot f-(\operatorname{div}f)\operatorname{Id})$, it suffices to show the estimate with $\mathcal{R}$ replaced by $\Delta^{-1}\nabla$. By Schauder estimates,
@@
$\|F_{j}\|_{C^{n,\kappa}}$ $\lesssim|\xi|^{n}\Big(\sum_{i=0}^{n}A_{i+j,\kappa}e^{-|\xi|^{2}/4}+\sum_{i=m}^{n+m}A_{i+j,\kappa}\Big)$
@@
$\quad\qquad\times|\lambda\xi|^{n+\kappa}\Big(|\lambda\xi|^{-1}\sum_{i=0}^{m-1}A_{i,\kappa}e^{-|\lambda\xi|^{2}/4}+\sum_{i=m}^{n+2m}A_{i,\kappa}\Big)$
@@
$\quad\qquad\times\Big(\sum_{i=0}^{m-1}|\lambda\xi|^{n-i-1}\|a\|_{C^{i,\kappa}}e^{-|\lambda\xi|^{2}/4}+\sum_{i=m}^{n+2m}|\lambda\xi|^{n-i}\|a\|_{C^{i,\kappa}}\Big).$
After bounding $e^{-|\lambda\xi|^{2}/4}$ by $e^{-\lambda^{2}/4}$, we sum over $\xi\in\mathbb{Z}^{d}\setminus 0$, which yields the claim. ∎

### v2

@@ 7. Proof of main results
For any test function $\varphi\in\mathcal{D}_{T}$ and $t_{0}\in(T_{*},T)$, we have
@@
for all $p<\infty$. Furthermore, $u$ lies in the Koch–Tataru path space $X_{T-T_{*}}$ and is weak-* continuous in time into $BMO^{-1}$. When $d\geq 3$, we additionally have
@@
First we address the boundedness in $\dot{W}^{-1,\infty}$ when $d\geq 3$. Recall that, by definition, $u=U+v^{N_{0}}+w^{N_{0}}$. Due to the criticality of the norm, we may equivalently estimate $w$ which, by construction, lies in $B_{X}(0,\delta)$. From this, (5.18), and (5.16), it follows
$\|w(t)\|_{\dot{W}^{-1,\infty}}$ $\lesssim_{\varepsilon}\int_{0}^{t}\|w\otimes w+2(v+U^{1/N_{0}})\odot w+f+2U^{1/N_{0}}\odot v\|_{C^{\varepsilon}}ds$
$\lesssim\int_{0}^{t}(s^{-1+\alpha-\varepsilon}+s^{-\frac{1}{2}(1+\varepsilon)}+1)ds$
@@
$\|R_{k}(t)\|_{\infty}$ $\lesssim_{\varepsilon}\int_{0}^{t}\|\bar{v}_{k+1}\otimes\bar{v}_{k+1}\|_{C^{\varepsilon}}\lesssim\int_{0}^{t}\sum_{j}N_{j,k+1}^{2+\varepsilon}e^{-2N_{j,k+1}^{2}t}ds\lesssim tN_{J_{d},k+1}^{2+\varepsilon}.$
@@
Next, we address the weak-* continuity in time into $BMO^{-1}$. For every $t\in[-T_{*},T-T_{*}]\setminus\{0\}$, $u$ is a classical solution and continuity is obvious. At $t=0$, left continuity is clear because $u$ agrees with $U$ on $T_{*},0]$. (Note that left continuity is vacuous if $T_{*}=0$.) To show right continuity at $t=0$, it suffices to show $u|_{[0,T-T_{*}]}\in X_{T-T_{*}}$ due to the recent result [24] of Hedong Hou. To this end, we decompose
@@
$\sup_{t\in(0,T-T_{*}]}t^{\frac{1}{2}-\epsilon}\|f(t)\|_{L^{\infty}}<\infty\implies f\in X_{T-T_{*}}$ (7.1)
@@
$e^{t\Delta}N_{j,k}\Delta\psi_{j,k}+\bar{v}_{j,k}$ $=N_{j,k}\Delta(e^{t\Delta}-e^{-N_{j,k}^{2}t})\psi_{j,k}$
@@
That $u_{2}\in X_{T-T_{*}}$ is immediate from Proposition 5.2 and (7.1).
@@
where the supremum is taken over cubes $Q\subset\mathbb{R}^{d}$; clearly, by periodicity, it suffices to consider only cubes below side length $1$, say. We proceed as in [12]*Proposition 3.8. Let $k(Q)$ and $j(Q)$ be such that
@@
Thus, the upper bound $\|u(t)\|_{L^{\infty}}\lesssim\frac{1}{\sqrt{t}}$ on $[0,T)$. To obtain the lower bound, we note that there exists a sequence of times $t_{n}\to 0+$ such that the principal velocity field $v$ satisfies
@@
We claim that the family of solutions $u^{(\sigma_{n})}$ converges to $u^{(0)}$ in $L_{t}^{\infty}([0,T];BMO^{-1}_{x}(\mathbb{T}^{d}))$ with the weak-* $BMO^{-1}$ topology. Indeed, by the rescaling, $\hat{v}^{C/\sigma_{n}}+\hat{w}^{C/\sigma_{n}}$ is supported only in $(C/\sigma_{n})\mathbb{Z}^{d}$. Since they are zero average, the support must be contained in $\{\xi\in\mathbb{Z}^{d}:|\xi|>C/\sigma_{n}\}$. Further, by criticality, we have
@@
Having shown inclusion in the Koch–Tataru space on the lifetime of the solution, the uniform bound in $BMO^{-1}$ follows from the theorem of [24]. Thus we conclude that for any test function $\varphi$ in the predual, we have
@@
$\lesssim\|P_{>C/(2\sigma_{n})}\varphi\|_{L_{t}^{1}\dot{F}^{1}_{1,2}}\|u^{(\sigma_{n})}-U\|_{L_{t}^{\infty}BMO^{-1}}.$
@@
$\|\nabla^{n}e^{t\Delta}\mathcal{R}\mathbb{P}(ae^{i\xi\cdot x})\|_{C^{\kappa}}$ $\lesssim|\xi|^{n+\kappa}\left(|\xi|^{-1}\sum_{i=0}^{m-1}A_{i,\kappa}e^{-|\xi|^{2}t/4}+\sum_{i=m}^{n+2m}A_{i,\kappa}\right).$
@@
The Leray projection is well-known to be bounded on $C^{\kappa}$ so it can be removed immediately. For the rest, we recall the following identity from [35]*Prop. 5.2:
@@
Since we can write $\mathcal{R}f=\Delta^{-1}(2\nabla\odot f-(\operatorname{div}f)\operatorname{Id})$, it suffices to show the estimate with $\mathcal{R}$ replaced by $\Delta^{-1}\nabla$. By Schauder estimates,
@@
$\|F_{j}\|_{C^{n,\kappa}}$ $\lesssim|\xi|^{n}\Big(\sum_{i=0}^{n}A_{i+j,\kappa}e^{-|\xi|^{2}t/4}+\sum_{i=m}^{n+m}A_{i+j,\kappa}\Big)$
@@
$\quad\qquad\times|\lambda\xi|^{n+\kappa}\Big(|\lambda\xi|^{-1}\sum_{i=0}^{m-1}A_{i,\kappa}e^{-|\lambda\xi|^{2}t/4}+\sum_{i=m}^{n+2m}A_{i,\kappa}\Big)$
@@
$\quad\qquad\times\Big(\sum_{i=0}^{m-1}|\lambda\xi|^{n-i-1}\|a\|_{C^{i,\kappa}}e^{-|\lambda\xi|^{2}t/4}+\sum_{i=m}^{n+2m}|\lambda\xi|^{n-i}\|a\|_{C^{i,\kappa}}\Big).$
After bounding $e^{-|\lambda\xi|^{2}t/4}$ by $e^{-\lambda^{2}t/4}$, we sum over $\xi\in\mathbb{Z}^{d}\setminus 0$, which yields the claim. ∎